Three years ago,Ali was eight times as old as Adi.Three years from now,Ali will only be five times as old as Adi.Find their present ages.
Found 2 solutions by math_helper, greenestamps:
Answer by math_helper(2461)



Let L = Ali's age now
and D = Adi's age now
From the 3yrs ago info:
(1) L-3 = 8(D-3) --> L = 8D-21
From the 3yrs from now info:
(2) L+3 = 5(D+3) --> L = 5D+12
L=L --> 8D-21 = 5D+12
Solve this to get D=11 (Adi is 11)
Plug D=11 into (1) or (2) and solve for L: L = 67 (Ali is 67)
Check:
67-3 = 64
8(11-3) = 64 (ok, using 3yrs ago info)
67+3 = 70
5(11+3) = 70 (ok, using 3yrs from now info)

Answer by greenestamps(13334)



The solution from the other tutor is fine....

Here is a different way to solve the problem with a different setup.

Given that Ali was 8 times as old as Adi 3 years ago...

Let x = Adi's age 3 years ago
Then 8x = Ali's age 3 years ago

Given that Ali will be 5 times as old as Adi 3 years from now (6 years after 3 years ago)...

x+6=Adi's age 3 years from now
8x+6=Ali's age 3 years from now

8x+6=5(x+6)
8x+6=5x+30
3x=24
x=8

Their ages 3 years ago were x=8 and 8x=64; their ages now are 11 and 67.

ANSWERS: Adi is now 11; Ali is now 67